1.1 L of a 2.50 M HNO3 solution is added to water so that the final volume is 6.0 L
Answer #1

Given

Initial volume = 1.1 L (V1)

Initial molarity = 2.50 M (M1)

Final volume = 6.0 L (V2)

I think so you require the final molarity (M2) as that is missing

We know that

M1V1 = M2V2

2.5 M * 1.1 L = M2 * 6.0 L

M2 = 2.5 M * 1.1 L / 6.0 L = 0.46 M